# On the fixed locus of the antisymplectic involution of an EPW cube

### v1

@@ Abstract.
Hyperkähler manifolds, antisymplectic involutions, Lagrangian fibrations, moduli spaces, Bridgeland stability, LLV decomposition.
@@ 1. Introduction
  • use this family and the degeneration methods of Flappan–Macrì–O’Grady–Saccà, who studied in [FMOS1] and [FMOS2] the fixed loci of antisymplectic involutions on hyper-Kähler varieties with a polarization of square $2$ to compute some numerical invariants of the fixed locus.
EPW cubes are defined as follows. Let $V_{6}$ be a 6-dimensional complex vector space and consider the symplectic space ${\bigwedge\hskip-2.56073pt^{3}}\hskip 0.56905pt{V}_{6}$. For any Lagrangian subspace $A\subseteq\textstyle{\bigwedge\hskip-2.56073pt^{3}}\hskip 0.56905pt{V}_{6}$, we consider the degeneracy loci
${\mathsf{Z}}^{\geq i}_{A}\coloneqq\{[U]\in\operatorname{\mathsf{Gr}}(3,V_{6})\mid\dim(A\cap\textstyle{\bigwedge\hskip-2.56073pt^{2}}\hskip 0.56905pt{U}\wedge V_{6})\geq i\}$
@@
When $A$ contains no decomposable vectors, that is, when the intersection ${\bf P}(A)\cap\operatorname{\mathsf{Gr}}(3,V_{6})\subseteq{\bf P}(\textstyle{\bigwedge\hskip-2.56073pt^{3}}\hskip 0.56905pt{V}_{6})$ is empty, the stratum ${\mathsf{Z}}_{A}\coloneqq{\mathsf{Z}}^{\geq 2}_{A}$ is a normal integral 6-fold and its singular locus is the normal integral 3-fold ${\mathsf{Z}}^{\geq 3}_{A}$. Moreover, there exists a double cover
@@
constructed in [DK, Theorem 5.7], branched along ${\mathsf{Z}}^{\geq 3}_{A}$. The variety $\widetilde{\mathsf{Z}}_{A}$ is called an EPW cube.
@@
$[{\mathsf{W}}_{A}]^{2}=\chi_{\mathrm{top}}({\mathsf{W}}_{A})=-1200.$
@@
As observed in Remark 3.2, the Lagrangian submanifold ${\mathsf{W}}_{A}$ of $\widetilde{\mathsf{Z}}_{A}$ is atomic in the sense of [Be, Definition 1.1]. After [Ma3, Section 7.4], a natural question is to ask whether it is rigid inside $\widetilde{\mathsf{Z}}_{A}$. In the second part of the paper (Sections 4 and 5), we use the degeneration method of [FMOS1] and [FMOS2] to show that the answer is positive.
@@
Let $\widetilde{\mathsf{Z}}_{A}$ be a smooth EPW cube with associated involution $\iota_{A}$. The fixed locus ${\mathsf{W}}_{A}$ of $\iota_{A}$ is a rigid atomic Lagrangian submanifold, namely its first Betti number is 00.
The degeneration we use is interesting in its own right and complements the study in [Ri] of another family of singular EPW cubes. It is constructed as follows. Let $(S,L)$ be a polarized K3 surface of degree $2$, with associated double cover $S\to{\bf P}^{2}$ branched over a smooth sextic curve $\Gamma\subseteq{\bf P}^{2}$. The class $2L-\delta$ is big and nef on the Hilbert scheme $S^{[3]}$, and it has square $4$ and divisibility $2$. Therefore, it induces a contraction $S^{[3]}\to X$, and the involution of $S$ given by the double cover induces an involution $\iota$ on $X$.
@@
Results from [BM1] on contractions of moduli spaces of sheaves on K3 surfaces allow us to explicitly describe $X$, the involution $\iota$, and its fixed locus. The latter is a disjoint union of two components, one of dimension $2$ (which does not deform), and $F_{3}$, of dimension $3$. The normalization of the component $F_{3}$ is a birational morphism
@@
As in [FMOS2], the most delicate part is showing that the (schematic) central fiber of $\mathscr{Y}\to B$ is equal to $F_{3}$, and in particular proving that it is reduced. This is done in Section 4.3 where, using the Kuranishi map, we determine the local structure the fixed locus of $\iota$.
@@
In Section 3, we specialize the previous analysis to the case of EPW cubes, to obtain Theorem 1.1. We also compute the cohomology class $[{\mathsf{W}}_{A}]$ (Proposition 3.1) and the Chern classes of ${\mathsf{W}}_{A}$ (Section 3.2).
@@ Acknowledgements
I would like to thank my advisor, Olivier Debarre, for suggesting me this problem and for his insights and hints. I am also grateful to Emanuele Macrì for the helpful discussions on the study of fixed loci in degenerations, and to Mirko Mauri for pointing out Kollár’s result [Ko, Corollary 2.64]. I thank Pietro Beri for useful discussions, and Francesco Denisi and Vanja Zuliani for their interest in this work.
[…]
@@
(the so-called Verbitsky component). Let $c_{2},c_{4},c_{6}$ be the nonzero Chern classes of $X$. One has ([Ma2, Lemma 1.5])
@@
The nonzero Chern numbers are ([EGL, Theorem 0.1])
@@
[…]
@@
where $\mathfrak{S}_{6-2k}$ is the group of permutations of $6-2k$ elements.
@@
Following [GKLR, Section 2], we denote by $\bar{V}$ the vector space $H^{2}(X,{\bf Q})$ with the (rational) Beauville–Bogomolov–Fujiki form $q$, and by $V$ the Mukai completion
@@
where $V_{(3)}$ is the Verbitsky component $\operatorname{SH}^{2\bullet}(X,{\bf Q})$ and $V_{(1,1)}$ is ${\bigwedge\hskip-2.56073pt^{2}}\hskip 0.56905pt{V}$. Furthermore, under the action of $\mathfrak{so}(\bar{V})$, the cohomology groups decompose further into the following representations
$\begin{array}[]{rcl}H^{0}(X,{\bf Q})&=&{\bf Q},\\ H^{2}(X,{\bf Q})&=&\bar{V},\\ H^{4}(X,{\bf Q})&=&\operatorname{Sym}^{2}\!\bar{V}\oplus\bar{V},\\ H^{6}(X,{\bf Q})&=&\operatorname{Sym}^{3}\!\bar{V}\oplus\textstyle{\bigwedge\hskip-2.56073pt^{2}}\hskip 0.56905pt{\bar{V}}\oplus{\bf Q}\eta,\\ H^{8}(X,{\bf Q})&=&\operatorname{Sym}^{2}\!\bar{V}\oplus\bar{V},\\ H^{10}(X,{\bf Q})&=&\bar{V},\\ H^{12}(X,{\bf Q})&=&{\bf Q},\end{array}$
where the summands in the fourth line are pairwise orthogonal with respect to the intersection pairing. [^3] The class $\eta$ was studied in [HHT, Proposition 6.1]. There are further decompositions
@@
into irreducible representations, whereas ${\bigwedge\hskip-2.56073pt^{2}}\hskip 0.56905pt{\bar{V}}=\bar{V}_{(1,1)}$ is already irreducible.
For $X$ very general (so nonprojective), the (special) Mumford–Tate algebra of $X$ is equal to the algebra $\mathfrak{so}(\bar{V})$ ([GKLR, Proposition 2.38]). Therefore the space of Hodge classes in $H^{\bullet}(X,{\bf Q})$ coincides with its $\mathfrak{so}(\bar{V})$-invariant part ([GKLR, Section 2.3]): it is given by
@@ 2.2. Action of automorphisms on the LLV algebra
By [F, Lemma 2.3], the application $h\mapsto(\varphi^{*})h(\varphi^{*})^{-1}$ induces an automorphism of $\mathfrak{g}$. In particular, $\mathfrak{g}\varphi^{*}(W)=\varphi^{*}\mathfrak{g}(W)=\varphi^{*}(W)$ for any $\mathfrak{g}$-invariant subspace $W$. ∎
@@
Since the map from the Verbitsky component to $H^{*}(X,{\bf Q})$ is given by cup product, one has $\varphi^{*}_{(3)}=\gamma_{(3)}$. For $\varphi^{*}_{(1,1)}$, the same argument as in [BGGG, Proposition 3.3] shows the following.
@@
We follow the proof of [BGGG, Proposition 3.3], borrowing its notation. Since $\varphi^{*}_{(1,1)}$ with $\gamma_{(1,1)}$ are both automorphisms of $V_{(1,1)}$, they satisfy
@@
Therefore $\varphi^{*}_{(1,1)}\circ\gamma_{(1,1)}^{n-1}$ is a $\mathfrak{g}$-isomorphism of $V_{(1,1)}$, hence, it is equal to $\lambda\cdot\operatorname{Id}$ by Schur’s lemma, for some $\lambda\in{\bf Q}$. Comparing orders, we obtain $\lambda^{n}=\operatorname{Id}$, hence $\lambda$ is $1$ for $n$ odd, and equal to ±$\pm 1$ for even $n$. ∎
@@
$\begin{array}[]{rcl}H^{0}(X,{\bf Q})&=&{\bf Q},\\ H^{2}(X,{\bf Q})&=&\bar{T}\oplus{\bf Q},\\ H^{4}(X,{\bf Q})&=&(\operatorname{Sym}^{2}\!\bar{T}\oplus\bar{T}\oplus{\bf Q})\oplus(\bar{T}\oplus{\bf Q}),\\ H^{6}(X,{\bf Q})&=&(\operatorname{Sym}^{3}\!\bar{T}\oplus\operatorname{Sym}^{2}\!\bar{T}\oplus\bar{T}\oplus{\bf Q})\oplus(\textstyle{\bigwedge\hskip-2.56073pt^{2}}\hskip 0.56905pt{\bar{T}}\oplus\bar{T})\oplus{\bf Q}.\\ \end{array}$
@@
The space $\operatorname{Hdg}^{6}(X,{\bf Q})$ also has dimension 3 and intersects the Verbitsky component in dimension $2$. Up to multiplication by a nonzero rational number, we obtain that $h\lambda$ is the absolute Hodge class $\eta$ from 8. By [HHT, Proposition 6.1], its square $4$ (modulo squares of nonzero rational numbers). The following lemma gives us a basis of $\mathrm{Hdg}^{\bullet}(X,{\bf Q})$.
@@
$\operatorname{Hdg}^{12}(X,{\bf Q})$ $={\bf Q}h^{6}.$
@@
[…]
@@
[…]
@@
$[W]\cdot h^{3}=\overline{[W]}\cdot h^{3}=b\Bigl{(}hc_{2}-\frac{12}{q(h)}h^{3}\Bigr{)}\cdot h^{3}=-72\,q(h)^{2}b,$

### v2

@@ Abstract.
Hyperkähler manifolds, antisymplectic involutions, Lagrangian subvariety, moduli spaces, Bridgeland stability, LLV decomposition.
@@ 1. Introduction
  • use this family and the degeneration methods of Flapan–Macrì–O’Grady–Saccà, who studied in [FMOS22] and [FMOS23] the fixed loci of antisymplectic involutions on hyper-Kähler varieties with a polarization of square 22 to compute some numerical invariants of the fixed locus.
EPW cubes are defined as follows. Let $V_{6}$ be a 6-dimensional complex vector space and consider the symplectic space ${\bigwedge\thinspace^{3}}\thinspace{V}_{6}$. For any Lagrangian subspace $A\subseteq\textstyle{\bigwedge\thinspace^{3}}\thinspace{V}_{6}$, we consider the degeneracy loci
${\mathsf{Z}}^{\geq i}_{A}\coloneqq\{[U]\in\operatorname{\mathsf{Gr}}(3,V_{6})\mid\dim(A\cap\textstyle{\bigwedge\thinspace^{2}}\thinspace{U}\wedge V_{6})\geq i\}$
@@
When $A$ contains no decomposable vectors, that is, when the intersection ${\bf P}(A)\cap\operatorname{\mathsf{Gr}}(3,V_{6})\subseteq{\bf P}(\textstyle{\bigwedge\thinspace^{3}}\thinspace{V}_{6})$ is empty, the stratum ${\mathsf{Z}}_{A}\coloneqq{\mathsf{Z}}^{\geq 2}_{A}$ is a normal integral 6-fold and its singular locus is the normal integral 3-fold ${\mathsf{Z}}^{\geq 3}_{A}$. Moreover, there exists a double cover
@@
constructed in [DK20, Theorem 5.7], branched along ${\mathsf{Z}}^{\geq 3}_{A}$. The variety $\widetilde{\mathsf{Z}}_{A}$ is called an EPW cube.
@@
${\mathsf{W}}_{A}]^{2}=\chi_{\mathrm{top}}({\mathsf{W}}_{A})=-1200.$
@@
As observed in Remark 3.2, the Lagrangian submanifold ${\mathsf{W}}_{A}$ of $\widetilde{\mathsf{Z}}_{A}$ is atomic in the sense of [Bec25, Definition 1.1]. After [Mar24, Section 7.4], a natural question is to ask whether it is rigid inside $\widetilde{\mathsf{Z}}_{A}$. In the second part of the paper (Sections 4 and 5), we use the degeneration method of [FMOS22] and [FMOS23] to show that the answer is positive.
@@
Let $\widetilde{\mathsf{Z}}_{A}$ be a smooth EPW cube with associated involution $\iota_{A}$. The fixed locus ${\mathsf{W}}_{A}$ of $\iota_{A}$ is a rigid atomic Lagrangian submanifold, namely its first Betti number is 0.
The degeneration we use is interesting in its own right and complements the study in [Riz24] of another family of singular EPW cubes. It is constructed as follows. Let $(S,L)$ be a polarized K3 surface of degree $2$, with associated double cover $S\to{\bf P}^{2}$ branched over a smooth sextic curve $\Gamma\subseteq{\bf P}^{2}$. The class $2L-\delta$ is big and nef on the Hilbert scheme $S^{[3]}$, and it has square $4$ and divisibility $2$. Therefore, it induces a contraction $S^{[3]}\to X$, and the involution of $S$ given by the double cover induces an involution $\iota$ on $X$.
@@
Results from [BM14a] on contractions of moduli spaces of sheaves on K3 surfaces allow us to explicitly describe $X$, the involution $\iota$, and its fixed locus. The latter is a disjoint union of two components, one of dimension $2$ (which does not deform), and $F_{3}$, of dimension $3$. The normalization of the component $F_{3}$ is a birational morphism
@@
As in [FMOS23], the most delicate part is showing that the (schematic) central fiber of $\mathscr{Y}\to B$ is equal to $F_{3}$, and in particular proving that it is reduced. This is done in Section 4.3 where, using the Kuranishi map, we determine the local structure the fixed locus of $\iota$.
@@
In Section 3, we specialize the previous analysis to the case of EPW cubes, to obtain Theorem 1.1. We also compute the cohomology class ${\mathsf{W}}_{A}]$ (Proposition 3.1) and the Chern classes of ${\mathsf{W}}_{A}$ (Section 3.2).
@@ Acknowledgements
I would like to thank my advisor, Olivier Debarre, for suggesting me this problem and for his insights and hints. I am also grateful to Emanuele Macrì for the helpful discussions on the study of fixed loci in degenerations, to Jieao Song for discussing with me Remark 2.7 and to Mirko Mauri for pointing out Kollár’s result [Kou93, Corollary 2.64]. I thank Pietro Beri for useful discussions, and Francesco Denisi, Giovanni Mongardi, and Vanja Zuliani for their interest in this work.
2. Hyper-Kähler manifolds of $\operatorname{{K3}}^{[3]}$-type
Let $X$ be a hyper-Kähler manifold of $\operatorname{{K3}}^{[3]}$-type. Its
nonzero Hodge numbers are
${b_{0}=1}$${1}$${b_{2}=23}$${1}$${21}$${1}$${b_{4}=299}$${1}$${22}$${253}$${22}$${1}$${b_{6}=2554}$  ${1}$${21}$${253}$${2004}$${253}$${21}$${1}$
@@ 2. Hyper-Kähler manifolds of $\operatorname{{K3}}^{[3]}$-type
(the so-called Verbitsky component). Let $c_{2},c_{4},c_{6}$ be the nonzero Chern classes of $X$. One has ([Mar20, Lemma 1.5])
@@
The nonzero Chern numbers are ([EGL01, Theorem 0.1])
@@
From [Son22, Lemma 2.1.7], we have [^1] All Chern numbers and Fujiki constants for hyper-Kähler manifolds of $\operatorname{{K3}}^{[n]}$-type can be computed using https://github.com/8d1h/bott.
@@
where $\mathfrak{S}_{6-2k}$ is the group of permutations of 62k6-2k elements.
@@
Following [GKLR22, Section 2], we denote by $\bar{V}$ the vector space $H^{2}(X,{\bf Q})$ with the (rational) Beauville–Bogomolov–Fujiki form $q$, and by $V$ the Mukai completion
@@
where $V_{(3)}$ is the Verbitsky component $\operatorname{SH}^{2\bullet}(X,{\bf Q})$ and $V_{(1,1)}$ is ${\bigwedge\thinspace^{2}}\thinspace{V}$. Furthermore, under the action of $\mathfrak{so}(\bar{V})$, the cohomology groups decompose further into the following representations
$\begin{array}[]{rcl}H^{0}(X,{\bf Q})&=&{\bf Q},\\ H^{2}(X,{\bf Q})&=&\bar{V},\\ H^{4}(X,{\bf Q})&=&\operatorname{Sym}^{2}\!\bar{V}\oplus\bar{V},\\ H^{6}(X,{\bf Q})&=&\operatorname{Sym}^{3}\!\bar{V}\oplus\textstyle{\bigwedge\thinspace^{2}}\thinspace{\bar{V}}\oplus{\bf Q}\eta,\\ H^{8}(X,{\bf Q})&=&\operatorname{Sym}^{2}\!\bar{V}\oplus\bar{V},\\ H^{10}(X,{\bf Q})&=&\bar{V},\\ H^{12}(X,{\bf Q})&=&{\bf Q},\end{array}$
where the summands in the fourth line are pairwise orthogonal with respect to the intersection pairing. [^3] The class $\eta$ was studied in [HHT12, Proposition 6.1]. There are further decompositions
@@
into irreducible representations, whereas ${\bigwedge\thinspace^{2}}\thinspace{\bar{V}}=\bar{V}_{(1,1)}$ is already irreducible.
For $X$ very general (so nonprojective), the (special) Mumford–Tate algebra of $X$ is equal to the algebra $\mathfrak{so}(\bar{V})$ ([GKLR22, Proposition 2.38]). Therefore the space of Hodge classes in $H^{\bullet}(X,{\bf Q})$ coincides with its $\mathfrak{so}(\bar{V})$-invariant part ([GKLR22, Section 2.3]): it is given by
@@ 2.2. Action of automorphisms on the LLV algebra
By [Flo22, Lemma 2.3], the application $h\mapsto(\varphi^{*})h(\varphi^{*})^{-1}$ induces an automorphism of $\mathfrak{g}$. In particular, $\mathfrak{g}\varphi^{*}(W)=\varphi^{*}\mathfrak{g}(W)=\varphi^{*}(W)$ for any $\mathfrak{g}$-invariant subspace $W$. ∎
@@
Since the map from the Verbitsky component to $H^{*}(X,{\bf Q})$ is given by cup product, one has $\varphi^{*}_{(3)}=\gamma_{(3)}$. For $\varphi^{*}_{(1,1)}$, the same argument as in [BGGG25, Proposition 3.3] shows the following.
@@
We follow the proof of [BGGG25, Proposition 3.3], borrowing its notation. Since $\varphi^{*}_{(1,1)}$ with $\gamma_{(1,1)}$ are both automorphisms of $V_{(1,1)}$, they satisfy
@@
Therefore $\varphi^{*}_{(1,1)}\circ\gamma_{(1,1)}^{n-1}$ is a $\mathfrak{g}$-isomorphism of $V_{(1,1)}$, hence, it is equal to $\lambda\cdot\operatorname{Id}$ by Schur’s lemma, for some $\lambda\in{\bf Q}$. Comparing orders, we obtain $\lambda^{n}=\operatorname{Id}$, hence $\lambda$ is $1$ for $n$ odd, and equal to $\pm 1$ for even $n$. ∎
@@
[…]
@@
The space $\operatorname{Hdg}^{6}(X,{\bf Q})$ also has dimension 3 and intersects the Verbitsky component in dimension 22. Up to multiplication by a nonzero rational number, we obtain that $h\lambda$ is the absolute Hodge class $\eta$ from (8). By [HHT12, Proposition 6.1], its square is $4$ (modulo squares of nonzero rational numbers).
The following lemma gives us a basis of $\mathrm{Hdg}^{\bullet}(X,{\bf Q})$.
@@
$\operatorname{Hdg}^{12}(X,{\bf Q})$ $={\bf Q}h^{6},$
where the class $\lambda$ depends on the polarization $h$, while the class $h\lambda$ is the absolute Hodge class $\eta$, and is orthogonal to the Verbitsky component $\operatorname{SH}^{3}(X,{\bf Q})$.
@@
Remark 2.5.
One also have the following relations involving $\lambda$. The class $\eta=h\lambda$ is orthogonal to the Verbitsky component $\operatorname{SH}^{3}(X,{\bf Q})$ and (up to multiply $\eta$ by a rational number), we can suppose $\eta^{2}=4$. Namely, in degree $12$ we have the following relations
$h^{2}\lambda^{2}=4,\qquad h^{4}\lambda=0,\qquad h^{2}c_{2}\lambda=0,\qquad c_{2}^{2}\lambda=0,$
where the last equality is obtain from the fact that $c_{2}^{2}$ is a linear combination of $h^{4}$ and $h^{2}c_{2}$.
Since $\operatorname{Hdg}^{5}(X)={\bf Q}\cdot h^{5}$, the above equalities together with Proposition 2.4 imply
$h\lambda^{2}=\frac{4}{15q(h)^{3}}h^{5},\qquad h^{3}\lambda=0,\qquad hc_{2}\lambda=0.$
We conclude this section by studying the class of Lagrangian submanifolds of hyper-Kähler manifolds of $\operatorname{{K3}}^{[3]}$}-type.
@@
Let $W$ be a Lagrangian submanifold of a very general polarized hyper-Kähler manifold $(X,h)$ of $\operatorname{{K3}}^{[3]}$}-type. The orthogonal projection to the Verbitsky component $\mathrm{SH}^{6}(X,{\bf Q})$ of the class $W]\in H^{6}(X,{\bf Q})$ is equal to
@@
$W]\cdot h^{3}=\overline{[W]}\cdot h^{3}=b\Bigl(hc_{2}-\frac{12}{q(h)}h^{3}\Bigr)\cdot h^{3}=-72\,q(h)^{2}b,$
@@
Remark 2.7.
Jieao Song pointed out to me that the above formula generalizes to any polarized hyper-Kähler variety $X$ of dimension $2n$ with polarization $h$, and any Lagrangian subvariety $F$ deforming with the polarization. In particular, following the same strategy of [Son22, Theorem 2.2.4], he shows that the orthogonal projection to the Verbitsky component $\operatorname{SH}^{2n}(X,{\bf Q})$ of the class $F]\in H^{2n}(X,{\bf Q})$ is equal to
(11) $\overline{[F]}=\frac{[F]\cdot h^{n}}{q(h)^{n}}\cdot\frac{\mu^{n}}{c_{X}}\overline{[\exp(h/\mu)\operatorname{td}_{X}^{1/2}]_{n}},\quad\text{where $\mu\coloneqq\sqrt{\frac{-q(h)}{2r_{X}}}$.}$
Here $c_{X}=\frac{C(1)}{(2n-1)!!}$ and $r_{X}\coloneqq\frac{C(c_{2})2^{n}n!(2n-1)}{(2n)!24C(1)}$ is the constant introduced in [Bec23, Equation (3.1)].
The key point is to generalize [Son22, Lemma 2.2.1] by showing that for any class $\alpha\in H^{2}(X,{\bf Q})$ there is an equality
$(\iota^{*}\alpha)^{n}\cdot[F]=\frac{[F]\cdot h^{n}}{q(h)^{n}}q(\alpha,L)^{n}.$
By writing $\operatorname{td}_{X}^{1/2}=\left(\operatorname{td}^{1/2}_{2k}\right)_{k}$ where $\operatorname{td}_{2k}^{1/2}\in H^{4k}(X,{\bf Q})$ and developing Equation (11), one obtain
(12) $\overline{[F]}=\frac{[F]\cdot h^{n}}{q(h)^{n}\cdot c_{X}}\sum_{2k\leq n}\left(\frac{-q(h)}{2r_{X}}\right)^{k}\frac{h^{n-2k}}{(n-2k)!}\ \overline{\operatorname{td}_{2k}^{1/2}}.$
[…]
$\overline{[F]}=\frac{[F]\cdot h^{3}}{q(h)^{3}}\left(\frac{h^{3}}{3!}-\frac{q(h)}{3\cdot 24}hc_{2}\right),$
which is exactly Proposition 2.6.